# Probing Visual Language Priors in VLMs

### v3

@@ 1 Introduction
Motivated by the results of ViLP, we propose Image-DPO, a self-improving approach to enhance VLM visual reasoning performance by increasing reliance on visual inputs. Our method employs self-generated VQAs using image generation and editing models (Podell et al., 2023; Ren et al., 2024; Brooks et al., 2023a) and applies controlled corruptions to create “good" and “bad" image pairs for DPO-like training (Rafailov et al., 2024). Experiments with open-source VLMs, including LLaVA-v1.5 (Liu et al., 2024a) and Cambrian (Tong et al., 2024), demonstrate its effectiveness. Moreover, we theoretically show that our objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022). The proposed ViLP dataset, Image-DPO code, and our synthetic data appear in https://huggingface.co/ViLP.
Figure 1: Sample data from ViLP. For the same question, ViLP  provides three distinct images and corresponding answers (upper-left corner). All questions follow a consistent structure, combining a distractor fact with a question. The Prior Answer (first column) can be directly inferred from the question, while Test Answers (second &\&& third column) rely on visual cues. Our answers are designed to be single words, and both the model and human evaluators are tasked with open-domain answering, rather than selecting from predefined options. To support this, we have developed a robust synonym and plural detection pipeline, ensuring that open-ended responses do not hinder the evaluation process. This approach also enables evaluation without relying on LLMs. Please refer to Appendix A.1 for more data samples from ViLP. Furthermore, we include both qualitative and quantitative comparison results with Winoground (Thrush et al., 2022), Whoops!(Bitton-Guetta et al., 2023a), and HallusionBench(Guan et al., 2023) in Appendix A.3.
@@ 2 Related Work
Vision Language Models and Language Priors: Multimodal reasoning is crucial for machine intelligence, with VLMs integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b, a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c, a; Zhao et al., 2023a; Chen et al., 2023; Zhu et al., 2024; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024). Inspired by the success of large language models (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a, b; Chiang et al., 2023) and pre-trained visual encoders (Radford et al., 2021; Desai & Johnson, 2020; Caron et al., 2021; Chen et al., 2024), many recent methods leverage relatively small vision-language paired datasets (Liu et al., 2024a; Tong et al., 2024) to fine-tune connectors between LLMs and visual backbones (Liu et al., 2024a). However, these datasets are far smaller than the vast text corpora for LLM pre-training (OpenAI, 2023b; Soldaini et al., 2024), and freezing the visual encoder and LLM parameters often preserves language biases, causing visual inputs to be overshadowed (Thrush et al., 2022; Sterz et al., 2024). This challenge is amplified by deliberately generated images that expose such biases, as shown in our study. Previous works and datasets (Goyal et al., 2016; Agrawal et al., 2017; Dancette et al., 2021; Wu et al., 2022; Ramakrishnan et al., 2018; Gouthaman & Mittal, 2020) addressed these issues with curated simulators (Johnson et al., 2016) or internet imagery (Zhang et al., 2016). In this paper, we present a novel VQA benchmark featuring carefully designed questions, fact-based distractors, rare-distribution answers, and image generation techniques to produce realistic visuals that challenge learned visual language priors (Figure 1).
@@
Self-Rewarding VLM: Self-rewarding LLM (Yuan et al., 2024) has shown that LLMs can generate and improving themselves in the process via Directed Peference Optimization (DPO) (Rafailov et al., 2024). This approach extends to VLMs by generating new answers for DPO training (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024b; Wang et al., 2024c, a; Yue et al., 2024; Liu et al., 2024b; Xiao et al., 2024). Our work aligns with these self-rewarding VLMs but differs in two key ways: (1) our proposed Image-DPO generates multiple images for a single question-image pair (rather than multiple answers); (2) rather than relying solely on existing images, Image-DPO pipeline creates diverse new images using pre-trained generative models (SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), InstructPix2Pix (Brooks et al., 2023a)). Furthermore, Image-DPO deliberately corrupts images to produce multiple degraded versions that serve as rejected data in DPO training. Concurrent works (Wang et al., 2024a; Xie et al., 2024) explore similar methodologies but lack a benchmark to verify enhanced visual focus, fail to establish theoretical connections between their proposed objective and DPO, and utilize limited image transformations (e.g., only randomly cropping). In contrast, we introduce ViLP (Section 3) to assess visual reasoning and provide theoretical foundations (Appendix B), alongside multi-category image corruptions (semantic editing, Gaussian blurring, pixelation).
@@ 3.1 Design Principles
To evaluate how VLMs handle learned visual language priors, we introduce ViLP, a specialized benchmark of out-of-distribution Question-Image-Answer (QIA) triplets guided by two core principles. First, Text-Only Inference ensures that each question can be answered with high confidence using textual clues alone. Second, Visual Influence requires that the correct answer—sometimes contradicting common sense—only emerges once an out-of-distribution image is considered. By forcing models to integrate both textual and visual information, ViLP reveals whether they truly engage in visual reasoning or merely rely on memorized patterns.
Mathematically, let $Q$ be a question, $I$ an image, and $A={a_{\text{prior}},\dots,a_{\text{test}},\dots}$ the set of possible answers. We define $P(a\mid Q)$ as the probability of answer $a$ given $Q$ alone and $P(a\mid Q,I)$ as the probability given both $Q$ and $I$. We consider a prior model $p$, which may represent either human cognition ($P_{\text{human}}$) or a VLM/LLM’s learned visual-language prior ($P_{\theta}$). For constructing our benchmark, we used the following guidances:
@@
In constructing ViLP, we use human cognition prior $P_{\text{human}}$ as guidance for its design. We evaluated the behavior of the learned priors of VLMs/LLMs by benchmarking $P_{\theta}$ against the human prior $P_{\text{human}}$.
@@ 3.2 Question-Image-Answer Generation
Following Criterion Three, $a_{\text{test}}$ should be highly improbable based on $Q$ alone yet the correct choice when paired with $I$. Since such images do not exist naturally, we use generative models like DALL·E-3 (Ramesh et al., 2021) and Flux to blend unusual elements that override typical language priors. We incorporate substantial human input and leverage advanced LLMs such as OpenAI-o1 and Claude-3.5-Sonnet to ensure alignment with all the criteria. More details are provided in Appendix A.2.
@@
Ultimately, we curated 300 questions, each paired with three distinct image-answer sets, totaling 900 QIA triplets. These cover a broad range from low-level recognition (texture, shape) to high-level reasoning (conceptual combinations, hallucinated components, proverbs). Table 1 summarizes their categorical distribution, with each question spanning an average of 1.6 categories. To reinforce text priors, we present a distractor fact before the question. Rigorous human review ensures that all final QIA triplets are clear and interpretable, as reflected by our human evaluation results in Table 2.
@@ 3.3 Dataset Evaluation
All of our questions are designed to be answered with a single word, as this design greatly simplifies evaluation protocol compared to assessing full sentences, without requiring LLMs for evaluation. We explicitly instruct the model to provide a single-word answer, and we evaluate the correctness of each response using a binary system. To ensure a fair evaluation, we devote significant effort to building a comprehensive set of synonyms and plural for each answer to detect other valid alternative answers. This ensures that the model is only penalized for actual errors, not for providing synonymous or alternative correct responses.
@@ 4.1 Objective
Figure 3: Qualitative examples. We show the results from GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B for some challenging cases.
@@
Formally, given an image $I_{w}$, a question $Q$, and its corresponding answer $A$, we generate a corrupted image $I_{l}$ via image-editing operations, including Gaussian blur, pixelation, or semantic modifications. The triplet $(Q,I_{l},A)$ forms a degraded question-image-answer pair compared to $(Q,I_{w},A)$. We train the model to distinguish between good and bad triplets using the objective below, where $\theta$ is the target VLM, $\theta_{\text{ref}}$ is the reference VLM (typically an earlier version of $\theta$), $S$ is the dataset of good and bad triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a scaling factor.
$L(\theta,\theta_{\text{ref}})$ $=-\mathbb{E}_{Q,I_{w},I_{l},A\sim S}\Bigl{[}$ (4)
$\quad\log\sigma\Bigl{(}\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}\;-\;\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})}\Bigr{)}\Bigr{]}$
@@ 5 Experiments
We introduce ViLP,  a benchmark of 300 questions, each with three unique images and corresponding answers (QIAprior and two QIAtest), yielding 900 QIAs. QIAprior aligns with common language priors (easily inferred from text alone), while QIAtest challenges these priors, requiring visual reasoning. ViLP features two evaluation settings: ViLPF (with distractor facts) and ViLPP (pure questions). We report average accuracy on QIAtest (Score) and on QIAprior (Prior) in Table 2.
@@
Is the QIA easy for humans? We begin by evaluating our benchmark through a human study. Participants achieved nearly perfect accuracy on ViLPF-Prior and over 98% on ViLPF-Score and ViLPP-Score, confirming that our question-image-answer combinations are unambiguous for human interpretation. Notably, despite QIAtest being designed as out-of-distribution examples, humans were still able to correctly distinguish them
@@
How do VLMs perform on our benchmark? Although our benchmark questions are distinguishable for humans, they are challenging for VLMs. Even the advanced VLM models like GPT-4o, have a clear performance gap (66.17% v.s. 98.33%) compared to humans’ performance on ViLPF-Score, indicating the difficulty of these questions for VLMs. Claude-3.5-Sonnet achieved the best score 70%, while most of the commercial VLMs are below 60%. Figure 3 highlights sample outputs from top commercial and open-source models, including GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B. They face significant challenges when addressing these cases in our ViLP, whereas humans can arrive at correct answers after consideration. Notably, it is encouraging to see that some open-source models achieved over 60% accuracy on ViLPF-Score, with performance nearing that of their commercial counterparts, including Llama-3.2-Vison and Molmo-72B.
@@
Comparison of ViLP with other VQA datasets: To highlight the distinctions between ViLP and existing benchmarks, including Winoground (Thrush et al., 2022), Whoops! (Bitton-Guetta et al., 2023a), and HallusionBench (Guan et al., 2023), we conduct a comparative analysis of both their high-level design principles and low-level data formats. This comparison incorporates qualitative and quantitative insights, offering a comprehensive understanding of their differences. Due to space constraints, detailed analyses are provided in Appendix A.3.
5.1 Image DPO
@@ 7 Acknowledgment
This work was supported by the LG AI Research grant. We also extend our gratitude to the OpenAI Researcher Access Program for providing the credits used to access OpenAI’s APIs. Finally, we greatly appreciate Chris Rockwell for his contributions to our human evaluation studies.
@@ A.2 More details in ViLPbenchmark data generation
A.3 Comparisons to other datasets

### v4

@@ 1 Introduction
Motivated by the results of ViLP, we propose Image-DPO, a self-improving approach to enhance VLM visual reasoning performance by increasing reliance on visual inputs. Our method employs self-generated VQAs using image generation and editing models (Podell et al., 2023; Ren et al., 2024; Brooks et al., 2023a) and applies controlled corruptions to create “good" and “bad" image pairs for DPO-like training (Rafailov et al., 2024). Experiments with open-source VLMs, including LLaVA-v1.5 (Liu et al., 2024a) and Cambrian (Tong et al., 2024), demonstrate its effectiveness. Moreover, we theoretically show that our objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022). The proposed ViLP dataset, Image-DPO code, and our synthetic data appear in https://vilp-team.github.io/.
Figure 1: Sample data from ViLP. For the same question, ViLP  provides three distinct images and corresponding answers (upper-left corner). All questions follow a consistent structure, combining a distractor fact with a question. The Prior Answer (first column) can be directly inferred from the question, while Test Answers (second &\&& third column) rely on visual cues. Our answers are designed to be single words, and both the model and human evaluators are tasked with open-domain answering, rather than selecting from predefined options. To support this, we have developed a robust synonym and plural detection pipeline, ensuring that open-ended responses do not hinder the evaluation process. This approach also enables evaluation without relying on LLMs. Please refer to Appendix A.1 for more data samples from ViLP. We investigate the impact of image styles in Appendix A.3, where we generate more realistic images using 4o image generation. Furthermore, we include both qualitative and quantitative comparison results with Winoground (Thrush et al., 2022), Whoops!(Bitton-Guetta et al., 2023a), and HallusionBench(Guan et al., 2023) in Appendix A.4.
@@ 2 Related Work
Vision Language Models and Language Priors: Multimodal reasoning is crucial for machine intelligence, with VLMs integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b, a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c, a; Zhao et al., 2023a; Chen et al., 2023; Zhu et al., 2024b; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024). Inspired by the success of large language models (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a, b; Chiang et al., 2023) and pre-trained visual encoders (Radford et al., 2021; Desai & Johnson, 2020; Caron et al., 2021; Chen et al., 2024), many recent methods leverage relatively small vision-language paired datasets (Liu et al., 2024a; Tong et al., 2024) to fine-tune connectors between LLMs and visual backbones (Liu et al., 2024a). However, these datasets are far smaller than the vast text corpora for LLM pre-training (OpenAI, 2023b; Soldaini et al., 2024), and freezing the visual encoder and LLM parameters often preserves language biases, causing visual inputs to be overshadowed (Thrush et al., 2022; Sterz et al., 2024). This challenge is amplified by deliberately generated images that expose such biases, as shown in our study. Previous works and datasets (Goyal et al., 2016; Agrawal et al., 2017; Dancette et al., 2021; Wu et al., 2022; Ramakrishnan et al., 2018; Gouthaman & Mittal, 2020) addressed these issues with curated simulators (Johnson et al., 2016) or internet imagery (Zhang et al., 2016). In this paper, we present a novel VQA benchmark featuring carefully designed questions, fact-based distractors, rare-distribution answers, and image generation techniques to produce realistic visuals that challenge learned visual language priors (Figure 1).
@@
Self-Rewarding VLM: Self-rewarding LLM (Yuan et al., 2024) has shown that LLMs can generate and improving themselves in the process via Directed Peference Optimization (DPO) (Rafailov et al., 2024). This approach extends to VLMs by generating new answers for DPO training (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024b; Wang et al., 2024c, a; Yue et al., 2024; Liu et al., 2024b; Xiao et al., 2024). Our work aligns with these self-rewarding VLMs but differs in two key ways: (1) our proposed Image-DPO generates multiple images for a single question-image pair (rather than multiple answers); (2) rather than relying solely on existing images (Zhu et al., 2024a), Image-DPO creates diverse new images using pre-trained generative models (SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), InstructPix2Pix (Brooks et al., 2023a)). Furthermore, Image-DPO deliberately corrupts images to produce multiple degraded versions that serve as rejected data in DPO training. Concurrent works (Wang et al., 2024a; Xie et al., 2024) explore similar methods but lack a benchmark to verify enhanced visual focus, fail to establish theoretical connections between their proposed objective and DPO, and utilize limited image transformations (e.g., only randomly cropping). In contrast, we introduce ViLP (Section 3) to assess visual reasoning and provide theoretical foundations (Appendix B), alongside multi-category image corruptions (semantic editing, Gaussian blurring, pixelation).
@@ 3.1 Design Principles
To evaluate how VLMs handle learned visual language priors, we introduce ViLP, a specialized benchmark of out-of-distribution Question-Image-Answer (QIA) triplets guided by two core principles. First, text-only inference ensures that each question can be answered with high confidence using textual clues alone. Second, visual inference requires that the correct answer—sometimes contradicting common sense—only emerges once an out-of-distribution image is considered. By forcing models to integrate both textual and visual information, ViLP reveals whether they truly engage in visual reasoning or merely rely on memorized patterns.
Mathematically, let $Q$ be a question, $I$ an image, and $A=\{a_{\text{prior}},\dots,a_{\text{test}},\dots\}$ the set of possible answers. We define $P(a\mid Q)$ as the probability of answer $a$ given $Q$ alone and $P(a\mid Q,I)$ as the probability given both $Q$ and $I$. We consider a prior model $p$, which may represent either human cognition ($P_{\text{human}}$) or a VLM/LLM’s learned visual-language prior ($P_{\theta}$). For constructing our benchmark, we used the following guidances:
@@
In designing ViLP, we leverage the human cognition prior $P_{\text{human}}$ as our guiding principle, ensuring each QIA configuration aligns with typical human expectations while requiring visual evidence to override strong textual assumptions. We then compare the learned priors of VLMs and LLMs, denoted $P_{\theta}$, $P_{\text{human}}$ to evaluate whether these models genuinely engage in visual reasoning rather than relying on memorized patterns.
@@ 3.2 Question-Image-Answer Generation
Following Criterion Three, $a_{\text{test}}$ should be highly improbable based on $Q$ alone yet the correct choice when paired with $I$. Since such images do not exist naturally, we use generative models like DALL·E-3 (Ramesh et al., 2021) and Flux to blend unusual elements that override typical language priors. We incorporate substantial human input and leverage advanced LLMs such as OpenAI-o1 and Claude-3.5-Sonnet to ensure alignment with all the criteria. More details, including text prompts and average cost, are provided in Appendix A.2. Note that as more advanced image generative models—such as the recently introduced 4o image generation—become available, we anticipate generating increasingly abundant and high-quality data for our benchmark, yet future updates will remain consistent with the dataset construction criteria outlined in Section 3.1.
@@
Ultimately, we curated 300 questions, each paired with three distinct image-answer sets, totaling 900 QIA triplets. These cover a broad range from low-level recognition (texture, shape) to high-level reasoning (conceptual combinations, hallucinated components, proverbs). Table 1 summarizes their categorical distribution, with each question spanning an average of 1.6 categories. To reinforce text priors, we present a distractor fact before the question. Rigorous human review ensures that all final QIA triplets are clear and interpretable, as reflected by our human evaluation results in Table 2. Besides, in Appendix A.3, we investigate the impact of image styles by generating more realistic images via 4o image generation and comparing them to those produced by DALL·E-3 (Ramesh et al., 2021) and Flux. We find that realistic images can increase the difficulty of the task, highlighting their importance for future studies.
@@ 3.3 Dataset Evaluation
All of our questions are designed to elicit single-word answers, an approach that is more efficient and more reliable than sentence-based evaluations that rely on LLM judgment. By avoiding using LLM, we reduce API fees, computational overhead, and the risk of occasional inaccuracies due to incorrect model reasoning. We explicitly instruct the model to provide a single-word answer, and we evaluate the correctness of each response using a binary system. To ensure a fair evaluation, we devote significant efforts to building a comprehensive set of synonyms and plural for each answer to detect other valid alternative answers. This ensures that the model is only penalized for actual errors, not for providing synonymous or alternative correct responses.
@@ 4.1 Objective
Figure 3: Qualitative examples. We show the results from GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B for some challenging cases. Please refer to Appendix A.6 for more failure case analysis.
@@
Formally, given an image $I_{w}$, a question $Q$, and its corresponding answer $A$, we generate a corrupted image $I_{l}$ via image-editing operations, including Gaussian blur, pixelation, or semantic modifications. The triplet $(Q,I_{l},A)$ forms a degraded question-image-answer pair compared to $(Q,I_{w},A)$. We train the model to distinguish between good and bad triplets using the objective 4, where $\pi_{\theta}$ is the target VLM, $\pi_{\text{ref}}$ is the reference VLM (typically an earlier version of $\pi_{\theta}$), $S$ is the dataset of good and bad triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a scaling factor.
$L(\pi_{\theta},\pi_{\text{ref}})$ $=-\mathbb{E}_{Q,I_{w},I_{l},A\sim S}\Bigl{[}$ (4)
$\quad\log\sigma\Bigl{(}\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{{\text{ref}}}(A\mid Q,I_{w})}\;-\;\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{{\text{ref}}}(A\mid Q,I_{l})}\Bigr{)}\Bigr{]}$
@@ 5 Experiments
We introduce ViLP, a new benchmark comprising 300 questions. Each question is paired with three unique images and their corresponding answers—one QIAprior and two QIAtest—for a total of 900 QIAs. The QIAprior examples (300 in total) align with common language priors (i.e., they can usually be answered correctly by relying on textual cues alone). In contrast, the QIAtest examples (600 in total) challenge these priors by requiring visual reasoning.
ViLP features two evaluation settings:
  • ViLPF, where both distractor facts and the questions are provided;
  • ViLPP, where only the questions themselves are given (i.e., no distractor facts).
We report two metrics in Table 2: average accuracy on QIAtest (noted as Score) and average accuracy on QIAprior (noted as Prior). Our benchmark emphasizes the performance in Score.
@@
Is the QIA easy for humans? We begin by evaluating our benchmark through a human study. Participants achieved nearly perfect accuracy on ViLPF-Prior and over 98% on ViLPF-Score and ViLPP-Score, confirming that our question-image-answer combinations are unambiguous for human interpretation. Notably, despite QIAtest being designed as out-of-distribution examples, humans were still able to correctly distinguish them.
@@
How do VLMs perform on our benchmark? Although our benchmark questions are distinguishable for humans, they are challenging for VLMs. Even the advanced VLM models like GPT-4o, have a clear performance gap (66.17% v.s. 98.33%) compared to humans’ performance on ViLPF-Score, indicating the difficulty of these questions for VLMs. Claude-3.5-Sonnet achieved the best score 70%, while most of the commercial VLMs are below 60%. Figure 3 highlights sample outputs from top commercial and open-source models, including GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B. They face significant challenges when addressing these cases in our ViLP, whereas humans can arrive at correct answers after consideration. Notably, it is encouraging to see that some open-source models achieved over 60% accuracy on ViLPF-Score, with performance nearing that of their commercial counterparts, including Llama-3.2-Vison and Molmo-72B. Additionally, we provide more detailed failure case analysis in Appendix A.6.
@@
Comparison of ViLP with other VQA datasets: To highlight the distinctions between ViLP and existing benchmarks, including Winoground (Thrush et al., 2022), Whoops! (Bitton-Guetta et al., 2023a), and HallusionBench (Guan et al., 2023), we conduct a comparative analysis of both their high-level design principles and low-level data formats. This comparison incorporates qualitative and quantitative insights as detailed in Appendix A.4.
@@ 7 Acknowledgment
This work was supported by the LG AI Research grant. We also extend our gratitude to the OpenAI Researcher Access Program for providing credits used to access OpenAI’s APIs. Additionally, we appreciate Chris Rockwell and Jiaming Yang for their contributions to our human evaluation studies, and thank Yongyi Yang for discussions regarding the proposition proof.
@@ A.2 More details in ViLPbenchmark data generation
Figure 10: Realistic image comparison. Each image pair shows our original benchmark data on the left and a corresponding realistic example generated by GPT-4o on the right.
A.3 Ablation studies: realistic images
Our benchmark data are currently generated by DALL·E-3 (Ramesh et al., 2021) and Flux, both of which produce cartoon-like, synthetic images rather than photorealistic ones. To assess the impact of image style, we regenerated a subset of 45 QIA pairs using GPT-4o’s latest image generation model to enhance realism, as illustrated in Figure 10. We then measured changes in model correctness when these more realistic images were used, with negative values indicating performance degradation. For definitions of these metrics, please refer to the beginning of Section 5. The results in Table 5 show that increased realism slightly reduces performance for the “Score” metric in most cases, while its effects on “Prior” are generally negligible. These ablation findings suggest that introducing more realistic images may increase the task’s difficulty, highlighting an important direction for future research.
$ViLP^{F}_{Prior}$ $ViLP^{F}_{Score}$ $ViLP^{P}_{Prior}$ $ViLP^{P}_{Score}$
GPT-4o 0 -2.2% 0 0
GPT-4o-mini 0 -1.1% 0 -2.2%
Claude-Sonnet-3.5 -2.2% -4.4% 0 -3.3%
Claude-Opus-3 2.2% -1.1% -2.2% -1.1%
Table 5: Impacts of Realistic Styles. Each value represents the change in correctness when replacing the original images with realistic ones (Realistic - Original). Negative values indicate a drop in performance, suggesting increased task difficulty. The metric definitions are provided in the beginning of Section 5.
A.4 Comparisons to other datasets